OpenStudy (anonymous):

Find tan if sec= sqrt26/5

OpenStudy (anonymous):

u mean \[\huge \frac{ \sqrt{26} }{ 5 }\] ?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

\[Ans \] \[\frac{ 1 }{ 5 }\]

OpenStudy (anonymous):

These are the choices. 5 -5

OpenStudy (anonymous):

Oh sry along with -sqrt25 and -1/5

OpenStudy (anonymous):

then it is -1/5

OpenStudy (anonymous):

Ok thanks again!